QuestionCoveBot:

|(-2/3F)3E ÷ (1/2F)2E| E means Exponent to left. F means Fraction.

TheSmartOne:

\(\left(\frac{a}{b}\right)^c = \frac{a^c}{b^c}\) First let's simplify the fractions and get rid of the exponents Remember a^3 means you multiply a by itself three times a*a*a

TheSmartOne:

Then when you divide the fractions remember a/b ÷ c/d to make that division into multiplication, we need to take the inverse of tbr second fraction so a/b * d/c